### v1

# Resilience of the autocatalytic feedback loop for gene regulation

Abstract
Gene expression in response to stimuli is regulated by transcription factors (TFs) through feedback loop motifs, aimed at maintaining the desired TF concentration despite uncertainties and perturbations. In this work, we consider a stochastic model of the autocatalytic feedback loop for gene regulation and we probabilistically quantify its resilience, i.e., its ability to preserve the equilibrium associated with a prescribed concentration of TFs, and the corresponding basin of attraction, in the presence of noise. We show that the formation of larger oligomers, corresponding to larger Hill coefficients of the regulation function, and thus to sharper non-linearities, improves the system resilience, even close to critical concentrations of TFs. Our formal results within a stochastic formulation relying on the Fokker-Planck equation are accompanied by numerical simulations.
1 Introduction and Motivation
Regulating gene expression in response to stimuli is essential for protein production and cell survival. Assessing the resilience of regulation pathways, i.e., their ability to withstand random perturbations and preserve crucial cell activities to survive and thrive, sheds light onto key biological mechanisms that sustain life and evolution, and enables the design of biomolecular circuits in synthetic biology.
Gene expression pathways, which can be successfully modelled using network motifs and control loops among responsive elements [1, 2], are fundamental to process external and internal stimuli and accordingly regulate crucial cell functions, such as enzymatic activity [3] or changes in gene expression during cell fate decisions [4]; they are also essential building blocks for circuits in synthetic biology [5].
The autocatalytic feedback loop motif, where a transcription factor (TF) acts as a positive regulator of its own production, is a simple and powerful model of gene expression regulation that captures essential mechanisms observed in living cells [6] to guarantee robust adaptation and homeostasis [7]. The model can exhibit multistability [8]: admitting multiple stable equilibria is evolutionarily advantageous, as it allows to rapidly switch between concentrations of TFs, thereby responding rapidly and reliably to changes in environmental or physiological conditions. For instance, the system modelling the expression of $\beta$-galactosidase in E. coli is bistable [9]: when lactose becomes available in the environment, the response pathway drives the concentration of the lac operon genetic inducer up to a critical threshold, yielding a sudden transition from low (“off” state, which saves energy) to high (“on”) concentrations of the $\beta$-galactosidase enzyme, which helps metabolise the nutrient and allows to feast on lactose before other bacteria, thus enabling survival. Then, as soon as nutrients deplete, another transition brings the system back to the “off” state, to avoid wasting resources. To efficiently manage gene translation and exploit nutrients to the fullest, transitions between stable equilibria must be tightly controlled and little sensitive to random deviations, environmental disturbances and intrinsic noise [10] to which cells are subject due e.g. to cell-to-cell heterogeneity [11], randomness in transport and binding of TFs [12], and transcriptional noise [13]. Understanding the evolutionary strategies and mechanisms that improve the resilience of the system in relation to its “off” and “on” states helps unveil how cells thrive in uncertain environments and inform the development of synthetic pathways.
We consider a model for autocatalytic feedback control of transcription and we investigate its resilience, formally defined as a probabilistic quantification of its ability to preserve a prescribed attractor and its corresponding basin of attraction in spite of stochastic noise [14], for its asymptotically stable equilibria. In particular, we assess the equilibria of the nominal deterministic system and their stability properties: the system is bistable for suitable parameter ranges and undergoes fold bifurcations. Then, we employ a stochastic approach to quantify its resilience in the presence of additive Gaussian white noise. Our analysis identifies factors that help the system cope with stochastic perturbations and avoid undesired random transitions between stable equilibria, thus approximately maintaining the desired TF concentrations, and illustrates how cells can thrive in noisy conditions.
Figure 1: Genetic regulation with autocatalytic feedback loop [15, 16]. The transcription factor TF-A, produced by gene tf-a, forms an oligomer ($n=3$) that, when phosphorylated (P), increases the transcription rate of tf-a, by promoting responsive DNA sequences. TF-A degrades at a constant rate.
2 Autocatalytic Feedback Loop: the Model
The autocatalytic feedback loop model for gene regulation introduced in [15] is visualised in Fig. 1. The single activator TF-A belongs to a pathway mediating cellular responses to stimuli and autoregulates its own transcription [17]. If the concentration $y\in[0,\infty)$ of TF-A is negligible, transcription of the tf-a gene occurs at a basal rate $r_{a}>0$. TF-A degrades following first-order kinetics, at a constant rate $\zeta>0$. Also, TF-A can form oligomers having concentration $y^{n}$, where $n\in\mathbb{N}$ represents how many monomers form the oligomer, which bind to responsive elements and, when phosphorylated, promote the transcription of tf-a; phosphorylation can be further regulated by external signals. We assume that binding processes are relatively rapid and close to equilibrium, so that the resulting increase in the transcription rate is captured by a monotonically increasing Hill function, with Hill coefficient $n$ and dissociation constant $K>0$ of oligomers from the responsive elements, which saturates to a maximal rate $\alpha>0$. All parameters are dimensionless, and the model does not explicitly consider the translation of mRNA into proteins. In a deterministic setting, the resulting ordinary differential equation (ODE) is $\dot{y}(\tau)=\alpha\frac{y(\tau)^{n}}{K+y(\tau)^{n}}-\zeta y(\tau)+r_{a}$. Considering the new variable $x\doteq y/K^{\frac{1}{n}}\in[0,\infty)$ and rescaling time as $t\doteq\tau/\zeta$ allows us to rewrite the ODE as
$\dot{x}(t)=f(x(t))=a\frac{x(t)^{n}}{1+x(t)^{n}}-x(t)+r,$ (1)
where $a=\alpha/(\zeta K^{\frac{1}{n}})>0$, $r=r_{a}/(\zeta K^{\frac{1}{n}})>0$, and the saturating Hill function $\frac{x^{n}}{1+x^{n}}$, for $n\to\infty$, converges to the Heaviside step function $\Theta(x-1)$.
We assess the system resilience through the lens of the rigorous formal definitions introduced in [14] for a family of ODE systems consisting of stochastic perturbations of a nominal deterministic system, aimed at probabilistically quantifying its ability to preserve a prescribed attractor $A$ (e.g., an asymptotically stable equilibrium) and the corresponding basin of attraction $B(A)$ in spite of noise. Our system family is =$\mathcal{F}=\{G_{\lambda}\}_{\lambda\in\mathcal{I}}$, with $\mathcal{I}=[0,\hat{\lambda})$ and $\hat{\lambda}>0$, where (1) is the nominal system $G_{\lambda_{0}}$, with $\lambda_{0}=0$, whereas the generic system $G_{\lambda}\in\mathcal{F}$ is described by the stochastic differential equation (SDE)
$\dot{x}(t)=f(x(t))+\lambda\eta(t)=a\frac{x(t)^{n}}{1+x(t)^{n}}-x(t)+r+\lambda\eta(t),$ (2)
where $\eta(t)$ is uncorrelated white noise with mean $\langle\eta\rangle=0$, variance $\sigma_{\eta}^{2}=1$ and intensity $\lambda\in\mathcal{I}$. As observed in [16], oligomers that require cooperative binding of more TF-A monomers, leading to a larger Hill coefficient $n$, allow for a better suppression of noise propagation close to the system equilibria. However, no formal results are available for the resilience of system (2) in relation to its asymptotically stable equilibria, which we investigate in this work.
3 Bistability of the Deterministic Model
Figure 2: (a, b) The intersections of functions $f_{1}(x)=ax^{n}/(1+x^{n})$ and $f_{2}(x)=x-r$ can be either one, two or three, and correspond to the equilibria of system (1); $\dot{x}=f(x)=f_{1}(x)-f_{2}(x)$ is positive when $f_{1}(x)>f_{2}(x)$, zero when $f_{1}(x)=f_{2}(x)$, and negative when $f_{1}(x)<f_{2}(x)$. (a) $f_{1}$ with $n=4$ and $a=1.5$; $f_{2}$ with $r\in\{0.1,0.186,0.35,0.428,0.6\}$. The intersections are two when $f_{2}$ is tangent to $f_{1}$, for $r=r_{c,1}\approx 0.186$ and $r=r_{c,2}\approx 0.428$, and three for $r_{c,1}<r<r_{c,2}$. The intersection is unique for $0<r<r_{c,1}$ and $r>r_{c,2}$. (b) $f_{1}$ with $n=4$ and $a\in\{1,1.6,3,105.5,600\}$; $f_{2}$ with $r=0.1$. The intersections are two when $f_{2}$ is tangent to $f_{1}$, for $a=a_{c,1}\approx 1.6$ and $a=a_{c,2}\approx 105.5$, and three for $a_{c,1}<a<a_{c,2}$. The intersection is unique for $0<a<a_{c,1}$ and $a>a_{c,2}$. (c, d) Zeros of $f(x)$ with $r=0.1$ for different values of $n$ and of $a$. (c) For $a=a_{c,1}(n)$, $f(x)$ is tangent to the $x$-axis at some point $\bar{x}(a_{c,1}(n))$: a fold bifurcation occurs that causes a transition from two to three equilibria. Solid lines: $n=4$, with $a_{c,1}(n)\approx 1.6$. Dashed lines: $n=7$, with $a_{c,1}(n)\approx 1.4$. (d) With $n=7$, for $a=a_{c,2}\approx 105.5$, $f(x)$ is tangent to the $x$-axis at some point $\bar{x}(a_{c,2}(n))$: a fold bifurcation occurs that causes a transition from three to two equilibria.
To analyse the qualitative behaviour of (1), we define functions $f_{1}(x)\doteq a\frac{x^{n}}{1+x^{n}}$ and $f_{2}(x)\doteq x-r$, such that $f(x)=f_{1}(x)-f_{2}(x)$. We consider $x\in[0,a+r]$, which is an invariant set for (1). In fact, (1) is a positive system ($x(0)\geq 0$ implies $x(t)\geq 0$ for all $t>0$, since $\dot{x}>0$ when $x=0$) and $\dot{x}<0$ when $x\geq a+r$, because $\sup_{x\in[0,\infty)}f_{1}(x)=a$.
The equilibria of system (1) are the intersections of the Hill function $f_{1}(x)$ and the line $f_{2}(x)$, as shown in Figs. 2a,b.
If $n=1$, $f_{1}$ is a Michaelis-Menten function, and this guarantees structural stability [18]: the system admits a unique equilibrium $\bar{x}_{1}=\frac{a+r-1+\sqrt{(a+r-1)^{2}+4r}}{2}$, which is structurally globally asymptotically stable, because $\dot{x}>0$ for $0\leq x<\bar{x}_{1}$ and $\dot{x}<0$ for $x>\bar{x}_{1}$, for all possible choices of $a>0$ and $r>0$. We next consider $n\geq 2$. Figs. 2c,d show $f(x)$ for a fixed value of $r$ and different values of $n\geq 2$ and $a$; the system equilibria, corresponding to the zeros of $f(x)$, can be either one (globally asymptotically stable), or two (one asymptotically stable and one unstable), or three (two asymptotically stable and one unstable); their stability can be assessed based on the sign of $f(x)$ in the intervals delimited by the equilibria.
Proposition 1
[…]
Proof 1
Computing the equilibrium values $\bar{x}_{i}$ for which $f(\bar{x}_{i})=0$ requires finding the roots of the polynomial $p(x)=-x^{n+1}+x^{n}(a+r)-x+r$, which in general is not possible analytically. In view of Descartes’ rule of signs, since there are three sign changes between consecutive nonzero coefficients, the polynomial has either one or three positive roots, counted with their multiplicity. This corresponds to three alternative scenarios: one equilibrium, two equilibria (of which one is associated with two coincident roots of $p(x)$), or three equilibria. As exemplified in Figs. 2a,b, which show the intersections of $f_{1}$ and $f_{2}$ for various parameter choices, all the scenarios are possible. Once all other parameters are fixed, the values $r_{c,1}<r_{c,2}$ of $r$ (or equivalently the values $a_{c,1}<a_{c,2}$ of $a$) for which $f_{2}$ is tangent to $f_{1}$ can be found by solving the nonlinear system
$\begin{cases}f_{1}(x)=f_{2}(x)\\ f_{1}^{\prime}(x)=f_{2}^{\prime}(x)\end{cases}$ (3)
Fig. 2a shows $f_{1}$ for given $n$ and $a$, and $f_{2}$ for varying values of $r$. If $0<r<r_{c,1}$ or $r>r_{c,2}$, there is a single equilibrium $\bar{x}_{1}>0$, which is globally asymptotically stable because $\dot{x}>0$ for $0\leq x<\bar{x}_{1}$ and $\dot{x}<0$ for $x>\bar{x}_{1}$. If $r=r_{c,1}$, there are two equilibria $0<\bar{x}_{1}<\bar{x}_{2}$; $\bar{x}_{1}$ is asymptotically stable with basin of attraction $[0,\bar{x}_{2})$ and $\bar{x}_{2}$ is unstable, since $\dot{x}>0$ for $0\leq x<\bar{x}_{1}$, while $\dot{x}<0$ for $\bar{x}_{1}<x<\bar{x}_{2}$ and $x>\bar{x}_{2}$. If $r_{c,1}<r<r_{c,2}$, there are three equilibria $0<\bar{x}_{1}<\bar{x}_{2}<\bar{x}_{3}$, where $\bar{x}_{1}$ and $\bar{x}_{3}$ are asymptotically stable, with basins of attraction $[0,\bar{x}_{2})$ and $(\bar{x}_{2},\infty)$ respectively, while $\bar{x}_{2}$ is unstable; in fact, $\dot{x}>0$ for $0\leq x<\bar{x}_{1}$, $\dot{x}<0$ for $\bar{x}_{1}<x<\bar{x}_{2}$, $\dot{x}>0$ for $\bar{x}_{2}<x<\bar{x}_{3}$, and $\dot{x}<0$ for $x>\bar{x}_{3}$. If $r=r_{c,2}$, there are two equilibria $0<\bar{x}_{1}<\bar{x}_{2}$, where $\bar{x}_{1}$ is unstable and $\bar{x}_{2}$ is asymptotically stable with basin of attraction $(\bar{x}_{1},\infty)$, since $\dot{x}>0$ for $0\leq x<\bar{x}_{1}$ and $\bar{x}_{1}<x<\bar{x}_{2}$, while $\dot{x}<0$ for $x>\bar{x}_{2}$. The same conclusions, with $r$ replaced by $a$, can be drawn when $r$ is fixed and $a$ varies, as in Fig. 2b.
Proposition 1 suggests that a fold bifurcation [19, 20] occurs when the number of equilibria changes from two to three (one equilibrium splits into two) and from three to two (two equilibria collide and merge). We consider $a$ as the bifurcation parameter and replace $f(x)$ by $f(x,a)$.
Proposition 2
[…]
Proof 2
For fixed $n$ and $r$, consider the values $a_{c,1}$ and $a_{c,2}$ of $a$ that satisfy (3). As shown in the proof of Proposition 1, when $a=a_{c,1}$ (respectively, $a=a_{c,2}$) the system admits two equilibria, $0<\bar{x}_{1}(a_{c,1})<\bar{x}_{2}(a_{c,1})$ with $\bar{x}_{1}(a_{c,1})$ asymptotically stable and $\bar{x}_{2}(a_{c,1})$ unstable (respectively, $0<\bar{x}_{1}(a_{c,2})<\bar{x}_{2}(a_{c,2})$ with $\bar{x}_{1}(a_{c,2})$ unstable and $\bar{x}_{2}(a_{c,1})$ asymptotically stable). We show that a fold bifurcation occurs both at $(\bar{x}_{2}(a_{c,1}),a_{c,1})$ and at $(\bar{x}_{1}(a_{c,2}),a_{c,2})$. According to [19, Theorems 3.1 and 3.2], the conditions for a fold bifurcation are: (I) $f(\bar{x}_{i},a_{c,j})=0$; (II) $f_{x}(\bar{x}_{i},a_{c,j})=0$; (III) $f_{a}(\bar{x}_{i},a_{c,j})\neq 0$; (IV) $f_{xx}(\bar{x}_{i},a_{c,j})\neq 0$. In our case, (I) and (II) are satisfied by construction, since they correspond to the two conditions in (3). (III) is $\frac{x^{n}}{1+x^{n}}|_{x=\bar{x}_{i}(a_{c,j})}\neq 0$, which is true, because all equilibria are strictly positive. (IV) requires that $\bar{x}_{i}\neq 0$, which is true, and $n\bar{x}_{i}^{n}+\bar{x}_{i}^{n}+1-n\neq 0$. If the latter condition is violated, $\bar{x}_{i}^{n}=\frac{n-1}{n+1}$ and then condition (II), $f_{x}(\bar{x}_{i},a_{c,j})=\frac{a_{c,j}n\bar{x}_{i}^{n-1}}{(1+\bar{x}_{i}^{n})^{2}}-1=0$, yields $a_{c,j}=\frac{4n}{(n^{2}-1)}\sqrt[n]{\frac{n-1}{n+1}}$, which contradicts our assumption on $a_{c,j}$. Since all conditions are verified, the system can be mapped to the normal form $\dot{z}=\beta\pm z^{2}$ of a fold bifurcation around the critical points [19, Section 3.3].
Fig. 3 shows bifurcation diagrams $\bar{x}(a)$ for $\dot{x}=f(x,a)$ with varying $n$. The system equilibria are computed symbolically for $n=2$, numerically for larger values of $n$. Their local stability is assessed through stability of the linearisation.
The analysis of system (1) has shown that, for $n\geq 2$, it can exhibit bistability (and thus switch between “off” and “on” states in response to external stimuli, see Section 1), and undergoes fold bifurcations for values $a_{c,1}$ and $a_{c,2}$ of the bifurcation parameter $a$ that, for fixed $r$, depend on $n$: $a_{c,1}(n)$ and $a_{c,2}(n)$. This is a so-called “resilience profile” [21]. As shown in the proof of Proposition 1, when the equilibrium $\bar{x}_{1}>0$ is unique, it is asymptotically stable with basin of attraction $B(\{\bar{x}_{1}\})=[0,\infty)$; when there are two equilibria, the asymptotically stable one is $\bar{x}_{1}(a_{c,1})$, with $B(\{\bar{x}_{1}(a_{c,1})\})=[0,\bar{x}_{2}(a_{c,1}))$, when $a=a_{c,1}$ and $\bar{x}_{2}(a_{c,2})$, with $B(\{\bar{x}_{2}(a_{c,2})\})=(\bar{x}_{1}(a_{c,2}),\infty)$, when $a=a_{c,2}$; when there are three equilibria, $\bar{x}_{1}$ and $\bar{x}_{3}$ are asymptotically stable, with $B(\{\bar{x}_{1}\})=[0,\bar{x}_{2})$ and $B(\{\bar{x}_{3}\})=(\bar{x}_{2},\infty)$.
We now wish to quantify the system’s ability to cope with stochastic disturbances. When two equilibria are closer (e.g., in the bifurcation diagram in Fig. 3, $|\bar{x}_{3}(q_{1})-\bar{x}_{2}(q_{1})|$ decreases when $n$ increases and $|\bar{x}_{1}(q_{2})-\bar{x}_{2}(q_{2})|$ decreases when $n$ decreases, with $q_{i}=a-a_{c,i}$), noise-induced switches [22] are intuitively more likely to occur, thus inducing transitions from the original equilibrium to another [23]. To assess the likelihood of such transitions in the presence of noise, we resort to the practical resilience framework [14].
Figure 3: (a) Bifurcation diagram $\bar{x}(a)$ for $\dot{x}=f(x,a)$, with $r=0.1$, computed analytically for $n=2$ (solid and dashed black lines for stable and unstable equilibria) and numerically for $n=3,5,8$ (lines in shades of blue and orange for stable and unstable equilibria). For small values of $a$, decreasing $a$ enables a transition from the high to the low stable equilibrium; transitions from the low to the high stable equilibrium occur if $a$ increases, but when $n>2$ they require a huge increase in $a$, see (b,c). (b, c) Bifurcation diagrams for $n=3$ with $a_{c,2}\approx 15$ (b) and for $n=5$ with $a_{c,2}\approx 153$ (c). For $n=8$, $a_{c,2}>250$ (not shown).
4 Resilience of the Stochastic Model
Consider the system family =$\mathcal{F}=\{G_{\lambda}\}_{\lambda\in\mathcal{I}}$ described by the parametric stochastic system (2) with noise intensity $\lambda$. To quantify the probability that, despite the stochastic noise, the system trajectories emanating from an initial condition $x_{0}\in B(A)$ converge to a prescribed neighbourhood of $A$, where $A$ is an asymptotically stable equilibrium (attractor) of the nominal deterministic system $G_{\lambda_{0}}$ given by (1), and $B(A)$ is the corresponding basin of attraction, we rely on the concept of asymptotic practical resilience introduced in [14, Definition 3], which formalises heuristic notions of resilience employed in systems biology [24].
Definition 1
[14] Consider the system family =$\mathcal{F}=\left\{G_{\lambda}\right\}_{\lambda\in\mathcal{I}}$ and let $(A,B(A))$ be an attractor-basin pair corresponding to $G_{\lambda_{0}}$. Fix a distance $\delta\geq 0$ and a confidence level $\gamma\in(0,1]$. The system $G_{\lambda}$ is $(\gamma,\delta)$-asymptotically practically resilient if, for all $x_{0}\in B(A)$,
$\mathbb{P}_{\lambda}\left(\limsup_{t\to\infty}\operatorname{dist}\left({x}(t;x_{0},\eta_{\lambda}),A\right)\leq\delta\right)\geq\gamma,$ (4)
where ${x}(t;x_{0},\eta_{\lambda})$ is the solution to system $G_{\lambda}$ emanating from $x_{0}$. The family $\mathcal{F}$ is $(\gamma,\delta)$-asymptotically practically resilient if, for all $x_{0}\in B(A)$,
$\inf_{\lambda\in\mathcal{I}\setminus\left\{\lambda_{0}\right\}}\mathbb{P}_{\lambda}\left(\limsup_{t\to\infty}\operatorname{dist}\left({x}(t;x_{0},\eta_{\lambda}),A\right)\leq\delta\right)\geq\gamma.$
[…]
If $\delta=0$ and $\gamma=1$, Definition 1 is equivalent to probabilistic robustness of the property “$(A,B(A))$ is an attractor-basin pair almost surely for the family $\mathcal{F}$”, which requires
$\mathbb{P}_{\lambda}\left(\left\{\lim_{t\to\infty}\operatorname{dist}(x(t;x_{0},\eta_{\lambda}),A)=0\right\}\right)=1$ for all $x_{0}\in B(A)$ and for all $\lambda\in\mathcal{I}$ [14].
The definition is based on the asymptotic properties of the system trajectories emanating from $x_{0}\in B(A)$ and requires them to converge to a $\delta$-neighbourhood of the attractor $A$ with probability at least $\gamma>0$. Allowing for $\delta>0$ is fundamental: due to the stochastic noise $\lambda\eta(t)$, the system trajectories emanating from $B(A)$ may not converge to $A$ (which is consistent with persistent noise-induced fluctuations observed in experiments), but the definition requires them to converge to a sufficiently small neighbourhood of $A$, with high probability, and not to alternative attractors (which could happen, e.g., due to a noise-induced switch). We can study how the probability of preserving a prescribed attractor-basin pair depends on the system parameters; by fixing $\gamma$, we can identify the system parameters that yield a desired confidence level.
4.1 Resilience quantification with the Fokker-Planck equation
The stochastic system (2) describes the evolution of a stochastic process, driven by white noise, having probability density function (PDF) $p(x,t)$. To assess the probability that a trajectory of (2), computed with the Itô formalism, converges to a prescribed set, we consider the corresponding Fokker-Planck equation (FPE), which describes the evolution of $p(x,t)$. In fact, any stochastic process whose PDF $p(x,t)$ satisfies the FPE associated with the SDE (2) is equivalent to the Itô solution of the SDE [25]. Following [25], a generic SDE in the Langevin equation form
$dx=a(x,t)dt+\sqrt{b(x,t)}\tilde{\eta}(t)dt,$
where the integral of the stochastic noise $\tilde{\eta}$ is a Wiener process, has an associated FPE for $p(x,t)$:
$p_{t}(x,t)=-[a(x,t)p(x,t)]_{x}+\frac{[b(x,t)p(x,t)]_{xx}}{2}=-J_{x}(x,t),$
where we have introduced the probability current
$J(x,t)=a(x,t)p(x,t)-\frac{1}{2}[b(x,t)p(x,t)]_{x}.$ (5)
The FPE associated with (2) is
$p_{t}(x,t)=-\left[f(x)p(x,t)\right]_{x}+\frac{\lambda^{2}}{2}p_{xx}(x,t),$ (6)
where $a(x,t)=f(x)$ and $\sqrt{b(x,t)}=\lambda$ do not depend on time. We set reflecting boundary conditions for $J(x,t)$ at $x=0$ and $x=+\infty$: $J(0,t)=J(+\infty,t)=0$, for all $t$.
We study the time-independent stationary solution $p_{s}(x)$ of (6), which is the limit of the solutions $p(x,t)$ for $t\to\infty$ in an appropriate metric, such as $L^{1}$ or relative entropy (see, e.g., [26, Sections 5.4 and 6.1] and [27] for more details). Hence, $p_{s}(x)$ captures the asymptotic PDF of the stochastic process $x$ solving (2) in the Itô sense. The probability in (4) of $x(t;x_{0},\eta_{\lambda})$ converging to a $\delta$-neighbourhood of the attractor $\bar{x}_{i}$, as per Definition 1, is thus
[…]
Following [25, Chapter 5.2.2], to obtain $p_{s}(x)$, we set $p_{t}(x,t)=0$ in (6). Since $J_{x}(x,t)=J_{x}(x)=0$ and the boundary conditions are zero, this amounts to setting $J=0$ in (5), with $a(x,t)=f(x)$ and $\sqrt{b(x,t)}=\lambda$, namely
$f(x)p_{s}(x)=\frac{\lambda^{2}}{2}p_{s}^{\prime}(x)=0,$ (8)
which yields
[…]
Since $f(x)$ is integrable, $p_{s}(x)$ can be rewritten as
$p_{s}(x)=\frac{\nu}{\lambda^{2}}\exp\left[\frac{1}{\lambda^{2}}\left(2rx-{x^{2}}+\frac{2ax^{n+1}\,\mathfrak{F}(x;n)}{n+1}\right)\right]\ ,$ (10)
where $\mathfrak{F}(x;n)={{}_{2}F_{1}}\left[1,\frac{n+1}{n};\frac{2n+1}{n};-x^{n}\right]$ is the Gaussian hypergeometric function [28, Chapter 15].
Therefore, $p_{s}(x)$ depends explicitly on $n$, $r$ and $a$. We study how it depends on the bifurcation parameter $a$ considered in Proposition 2, also in relation to the equilibria of the nominal deterministic system analysed in Proposition 1.
Proposition 3
[…]
Proof 3
Denoting by $h(x)$ the argument of the exponential in (9) and substituting the expression of $p_{s}(x)$ into (8) yields $p_{s}^{\prime}(x)=\frac{2}{\lambda^{2}}f(x)\frac{\nu}{\lambda^{2}}\exp[h(x)]=0$. Since the exponential function is always positive, this is equivalent to requiring $f(x)=0$, which is the equilibrium condition for system (1). Hence, the statement follows from the results in Section 3. The stationary points of $p_{s}(x)$ are attained for values of $x$ that are equilibria of (1). Considering the sign of $f(x)$ within the intervals of values of $x$ delimited by the zeros of $f(x)$, we have a maximum for $p_{s}(\bar{x}_{i})$ if $\bar{x}_{i}$ is asymptotically stable, a minimum for $p_{s}(\bar{x}_{i})$ if $\bar{x}_{i}$ is the unstable equilibrium in the three-equilibria case, and an inflection point for $p_{s}(\bar{x}_{i})$ if $\bar{x}_{i}$ is the unstable equilibrium in the two-equilibria case. Hence, $p_{s}(x)$ has two local maxima in the three-equilibrium case and a single maximum otherwise.
In view of Proposition 3, $p_{s}(x)$ is a good indicator of the system resilience in relation to the preservation of prescribed attractors: its maxima occur at the attractors, and the maxima are two when the deterministic system is bistable, because noise can drive the stochastic trajectories to a different basin of attraction, with a probability that increases with the noise intensity $\lambda$. Fig. 4 shows $p_{s}(x)$ for different values of ${n}$, for increasing noise intensity $\lambda$, for three different values of $a$, computed based on trajectories emanating from random, uniformly distributed initial conditions $x_{0}\in[0,\infty)$. For $a=0.8$ and $a=2.7$, system (1) is monostable for all considered $n$ and its equilibrium is at low values (“off”) and high values (“on”), respectively; for $a=a_{c,1}(n)+0.05$, the system is bistable. When $\lambda^{2}=0.01$, in the monostable cases, all the noisy trajectories converge very close to the attractor of the nominal deterministic system: $p_{s}(x)$ is an approximation of a Dirac delta centred at the attractor when $a=2.7$, and is centred at the attractor with a very narrow peak when $a=0.8$; in the bistable case ($a=a_{c,1}(n)+0.05$), $p_{s}(x)$ has a maximum at the low asymptotically stable equilibrium, a minimum at the unstable equilibrium and a second maximum at the high asymptotically stable equilibrium. When $\lambda^{2}=0.05$, the trajectories are more likely to converge to regions of the state space further away from the attractors. For all simulated noise intensities, when $n$ is larger, $p_{s}(x)$ has higher and narrower peaks centred at the equilibria, thus confirming the observation in [16] that noise suppression close to the equilibria improves when using more complex oligomers (composed of a larger number $n$ of monomers).
Figure 4: Computation of $p_{s}(x)$ from (10), for $r=0.1$ and $n\in\{2,3,5,8\}$. Columns: $\lambda^{2}=0.01$ (left) and $\lambda^{2}=0.05$ (right). Rows: $a=2.7$ (top); $a=a_{c,1}(n)+0.05$ (middle); $a=0.8$ (bottom). The constant $\nu$ is estimated numerically. The inset on the second row for $\lambda^{2}=0.01$ shows the second maximum for n¿2.
4.2 Numerical simulations
We can quantify $\mathcal{P}_{\lambda}$ numerically for an attractor $\bar{x}_{i}$, i.e., an asymptotically stable equilibrium of (1). To this aim, we generate random trajectories $x(t;x_{0},\eta_{\lambda})$ with $x_{0}\in B\left(\left\{\bar{x}_{i}\right\}\right)$, and assess how frequently they converge to $(\bar{x}_{i}-\delta,\bar{x}_{i}+\delta)$.
In particular, we perform numerical simulations of the system (2) using an Euler-Maruyama scheme, with $r=0.1$ and $a(n)=a_{c,1}(n)+0.05$, which guarantees bistability for all the considered choices of $n\in\{2,3,5,8\}$ (cf. Fig. 3), and with uniformly spaced initial conditions $x_{0}\in B(\{\bar{x}_{3}(n)\})=(\bar{x}_{2}(n),r+a(n)]$, where $r+a(n)$ is the upper limit of the invariant set for system (1) identified in Section 3.
For the various choices of $n$ and for a range of noise intensities $\lambda$, we compute the fraction of trajectories of (2) that have converged to $(\bar{x}_{3}(n)-\delta,\bar{x}_{3}(n)+\delta)$ at the time $\hat{t}=100s$ such that the trajectories of the deterministic system (1) have already reached their steady state. We set $\delta=0.25$, such that $|\bar{x}_{3}(n)-\bar{x}_{2}(n)|<\delta$ and $|r+a(n)-\bar{x}_{3}(n)|<\delta$ for all considered values of $n$. Fig. 5a shows $\mathcal{P}_{\lambda}$ depending on $n$ and $\lambda$: the probability of converging to a $\delta$-neighbourhood of the attractor is larger if $\lambda$ is smaller (as expected, since $\lambda$ is the noise intensity). If $n$ increases, $\mathcal{P}_{\lambda}$ first increases and then reduces: more complex oligomers (formed by a larger number $n$ of monomers) initially increase the resilience to noise, but then bring the system to lower levels of asymptotic resilience, linked to the small peaks of $p_{s}(x)$ for $\bar{x}_{3}(n)$, observed in Fig. 4.
For a fixed $\lambda=0.01$, Fig. 5b shows $\mathcal{P}_{\lambda}$ for different values of $\delta\in(0,0.6]$, where a smaller $\delta$ corresponds to a stricter requirement. For smaller values of $n$, $\mathcal{P}_{\lambda}$ is higher for larger $\delta$; conversely, given a fixed probability $\gamma^{*}$, the smallest $\delta$-neighbourhood of $\bar{x}_{3}(n)$ to which convergence occurs with probability at least $\gamma^{*}$ is smaller when $n$ is larger (more complex oligomers), which is favourable to reduce variability in the concentration of TFs, thus confirming the qualitative observations in [16].
Figure 5: (a) Dependence of $\mathcal{P}_{\lambda}$ on $\lambda$ and $n$, for system (2) with $r=0.1$, $a=a_{c,1}(n)+0.05$ and a fixed $\delta=0.25$. (b) Dependence of $\mathcal{P}_{\lambda}$ on $\delta$ and $n$, for system (2) with $r=0.1$, $a=a_{c,1}(n)+0.05$ and a fixed $\lambda=0.01$.
5 Conclusion
We have investigated the resilience of the autocatalytic feedback loop motif for gene regulation, where the positive feedback is enforced by oligomers formed by $n$ monomers. We have studied the equilibria of the deterministic system and their stability, showing that it can exhibit both monostable and bistable behaviours depending on the parameter values, and proved that the system undergoes fold bifurcations when crucial parameters are varied. In a stochastic framework, we have relied on the Fokker-Planck equation to assess the asymptotic practical resilience of the system in the presence of noise, as defined in [14]. We have analysed how different parameters affect the resilience of the system, and focused, in particular, on the oligomer size $n$. Our analysis highlights which biological designs can better meet prescribed resilience requirements in probability, in terms of preservation of an attractor-basin pair given a maximal noise intensity. Our results also enable future studies aimed at choosing the optimal oligomer size $n$ to enforce resilience requirements, accounting for the cost associated with oligomer production and disposal.
Acknowledgments
This work was funded by the European Union through the ERC INSPIRE grant (project number 101076926). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the European Research Council Executive Agency can be held responsible for them.
Code
The code to reproduce the results is at
https://github.com/daniele-proverbio/GeneAutocatalysis.